A Quick (.1 seconds!) Proof of Gigoujeu’s Two-Circle Theorem

Shalosh B. EKHAD

À la mémoire de Michel Jérôme Dufrénoy

In 1960, Francois Némorin Gigoujeu famously won the grand prix (a library of 3000 volumes!) in the concours organized by the Société Générale de Crédit Instructionnel (see [1]), by solving the following challenging question.

“On donne deux circonférences OO′: d’un point A pris sur O, on mène des tangentes à O′; on joint les points de contact de ces tangentes: on mène la tangente en A à la circonférence O; on demande le lieu du point d’intersection de cette tangente avec la corde des contacts dans la circonférence O′”

By rescaling and rotating, we can let O′ be x^{2} + y^{2} = 1 and O be (x − a)^{2} + y^{2} = r^{2}.

Running the Maple code (by D. Zeilberger), available from http://www.math.rutgers.edu/~zeilberg/tokhniot/fng,

I immediately confirmed Gigoujeu’s theorem that states that the desired locus is

a^{2}x^{4}+2a(r^{2}−a^{2}−1)x^{3}+(a^{2}y^{2} + a^{4} + 4 a^{2} − 2 a^{2}r^{2} + (r^{2} − 1)^{2}) x^{2}+2a ((r^{2} − 1)y^{2} + r^{2} − a^{2} − 1) x

+ ((r^{2} − 1)^{2} − a^{2}r^{2}) y^{2} + a^{2} = 0 .

See a diagram here: http://www.math.rutgers.edu/~zeilberg/PG/Gigoujeu.html .

Reference

1. Jules Verne, “Paris au XX^{e} siècle”, Hachette, 1994 .